OpenStudy (anonymous):

two large numbers of the fibonacci sequene are f 50= 12586269025 and 51=20365011074 what is the approximate value of the quotient f61/f60

OpenStudy (anonymous):

a1.618 b0.618 c2.618 d2 e3.142

OpenStudy (campbell_st):

well find the ratio of \[\frac{f_{51}}{f_{50}} = \] you can check it the sequence is 1, 1, 2, 3, 5, 8, 13, 21, so what is the ratio of 21/13 is it close to 13/8 and then compare it to \[\frac{f_{51}}{f_{50}} = \]

OpenStudy (anonymous):

how do i find the ratio

OpenStudy (campbell_st):

its just a fraction 13/8 = 1.625 21/13 = 1.61538 so what is 12586269025/20365011074 I'd expect it's getting close to 1.608

OpenStudy (anonymous):

i got 0.618033988749895

OpenStudy (campbell_st):

oops should be the other way around 20365011074/12586269025

OpenStudy (anonymous):

yep thnx so the answer should be B

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

did you recalculate it...?

OpenStudy (anonymous):

recalculate 20365011074/12586269025 yep

OpenStudy (campbell_st):

as the recalculation gives 1.61803 so which is a good choice...

OpenStudy (anonymous):

so im confused would it be A or B

OpenStudy (anonymous):

A